OpenStudy (anonymous):

please show me how to do this? (6b^3+5b-8)-(8b-1+4b^3)

OpenStudy (anonymous):

Result is\[2b ^{3} -3b -7\] The whole problem is addition and subtraction, start by removing parenthesis, to do this, distribute that minus sign through the second set of parentheses so it becomes -8b+1-4b^3 then eliminate the rest with addition and subtraction 6b^3 - 4b^3 = 2b^3 5b - 8b = -3b -8 + 1 = -7